# Extracting Essential and Disentangled Knowledge for Recommendation Enhancement

### v1

@@ Abstract.
Recommender models play a vital role in various industrial scenarios, while often faced with the catastrophic forgetting problem caused by the fast shifting data distribution, e.g., the evolving user interests, click signals fluctuation during sales promotions, etc. To alleviate this problem, a common approach is to reuse knowledge from the historical data. However, preserving the vast and fast-accumulating data is hard, which causes dramatic storage overhead. Memorizing old data through a parametric knowledge base is then proposed, which compresses the vast amount of raw data into model parameters. Despite the flexibility, how to improve the memorization and generalization capabilities of the parametric knowledge base is challenging. In this paper, we propose two constraints to extract Essential and Disentangled Knowledge from past data for rational and generalized recommendation enhancement, which improves the capabilities of the parametric knowledge base without increasing the size of it. The essential principle helps to compress the input into representative vectors that capture the task-relevant information and filter out the noisy information. The disentanglement principle reduces the redundancy of stored information and pushes the knowledge base to focus on capturing the disentangled invariant patterns. These two rules together promote rational compression of information for robust and generalized knowledge representations. For the parametric knowledge base, we train a knowledge extractor that extracts knowledge patterns of arbitrary order from past data and a knowledge encoder that memorizes the arbitrary order patterns, which serves as the retrieval key generator and memory network respectively in the following knowledge reusing phase. The whole process is regularized by the proposed two constraints. Extensive experiments on two datasets justify the effectiveness of the proposed method.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Figure 1. (a) The illustration of the overall process. The vast knowledge of old data is compressed into the parametric knowledge base. (b) The essential principle for compression, which encourages a compressed representation that captures the fundamental knowledge of data and filters out the noises. (c) The disentangled principle for compression, which reduces the redundancy of stored patterns and decomposes the invariance for better generalization.
@@
To solve the challenges above, we propose the Essential and Disentangled Knowledge base, where two constraints regularize a parametric knowledge base for robust and generalized knowledge preserving without increasing the size of the knowledge base. Concretely, the Essential principle aims to capture the task-relevant information (the red part in Figure 1.(b)) and filter out the noisy information (the grey part in Figure 1.(b)) in the data, which follows the Information Bottleneck principle (Tishby et al., 2000a) that finds the optimal trade-off between compression and prediction. By striking a balance between these two aspects, the principle enables the model to preserve sufficient information of data while keeping the representation minimal and invariant (Achille and Soatto, 2018), which contributes to lower storage cost and higher generalization capability. The Disentanglement principle eliminates redundant information among different patterns (the grey part in Figure 1.(c)) and decomposes the invariance of different patterns. This reduction of redundancy helps to improve memorization within limited model parameters. And at the same time, as depicted in (Montero et al., 2020), improving disentanglement may contribute to the generalization. The two rules together contribute to compressed and rational knowledge vectors that capture the fundamental knowledge of past data and have good generalization ability for future distribution.
@@
  • We propose a parametric knowledge base that compresses the massive old data documents into robust and generalized parametric knowledge, which is a modal-agnostic method that could offer diverse knowledge for various base models.
@@
  • The knowledge extraction and utilization of EDK is instance-level, which offers fine-grained information tailed for each instance for enhancement.
@@ 2. Related Work
2.1. Recommender Systems
This work is closely related to the research topics of recommender systems, which focus on capturing collaborative signals based on feature interaction learning and user behavior modeling.
@@ 2.1. Recommender Systems
For feature interaction learning, there are plenty of works mining the interactive patterns of categorical data. FM (Rendle, 2010) models features as low-dimensional embeddings and captures feature interactions by inner products. FFM (Juan et al., 2016) further develops field-aware interactions by multiple embeddings settings. NeuFM (He et al., 2017) proposes to use deep neural networks, improving FM by incorporating a multi-layer perceptron. Wide & Deep (Cheng et al., 2016) combines the strengths of linear models and DNNs for memorization and generalization respectively. DeepFM (Guo et al., 2017) utilizes an FM layer to replace the wide component in (Cheng et al., 2016) to model the pairwise feature interactions. xDeepFM (Lian et al., 2018) introduces the cross layer to replace the wide component and constructs limited high order feature interactions. PNN (Qu et al., 2018) proposes a product layer to better model the “AND” relations between features. DCN (Wang et al., 2017) learns both low-dimensional feature crossing and high-dimensional nonlinear features efficiently. AutoInt (Song et al., 2019) learns high-order feature interactions via the self-attention mechanism. AutoFIS (Liu et al., 2020) applies a gate for each feature interaction and selected limited order beneficial interaction by learning the gate values.
@@
User behavior modeling plays a crucial role in recommendation systems by providing valuable insights into user preferences, interests, and behaviors. By understanding how users interact with a system and make choices, recommendation algorithms can effectively personalize and tailor recommendations to individual users. DIN (Zhou et al., 2018) models user interests with attention mechanism. It reduces the influence of historical behavior information of commodities that are not related to the currently estimated advertisement on the current click estimation judgment. DIEN (Zhou et al., 2019) further points out that user interests evolve over time and modeled the evolving interests of users with the GRU module. DSIN (Feng et al., 2019) splits different sessions and better models the evolving property of user interests. MIMN (Pi et al., 2019) further categorizes user interests into different channels and models different aspects of user interests. SIM (Pi et al., 2020) proposes to model long-term user behaviors and aggregates them using the fast SimHash algorithm.
@@ 2.2. Model Invariance & Generalization
3. Problem Formulation
(6) $p(y|X,\theta),$
where $X=\left[x_{1},x_{2},\dots,x_{F}\right]$ is the input consisting of features from $F$ feature fields and $\theta$ is the model parameter. These methods make prediction based solely on target feature and model parameters. Another thread of methods that models user behavior patterns (Zhou et al., 2018, 2019; Pi et al., 2019; Feng et al., 2019; Pi et al., 2020) shows important impacts on making personalized recommendation and brings great performance gains, which can be formulated as
(7) $p(y|X,D_{his},\theta),$
(8) $KB_{\theta*}\leftarrow D_{old},$
(9) $p(y|X,KB_{\theta*},\theta),$
(10) $p(y|X,D_{his},KB_{\theta*},\theta),$
Figure 2. The framework of EDK . In the knowledge compression stage, we compress the essential and disentangled knowledge within old data into the parametric knowledge base. Concretely, we extract patterns within data instances with the knowledge extractor and memorize them through the knowledge encoder that could deal with inputs of arbitrary scales. The overall knowledge compression process is regularized by two principles: essential and disentangled for better generalization and robustness. During prediction, the target could access the frozen knowledge base for instance-wise knowledge and adapt the knowledge to inject it into arbitrary recommendation backbone for enhanced prediction.
@@ 4.1.1. Essential & Disentangled Principles
  • Essential. $\min I(X;\{\mathbf{s_{j}}\})-\alpha I(\{\mathbf{s_{j}}\};y)$. The generated knowledge representations should contain as much of the task-relevant information of the original input and as less of task-irrelevant information. It helps to distill complex data into a sufficient and minimal representation, enabling better generalization and robustness.
@@ 4.1.4. Objective
where $l_{ce}$ denotes the cross entropy loss of predicting labels on $\mathbf{c}$, $l_{reg}$ is the loss corresponding to the two principles proposed in Equation (13), and $l_{0}$ denotes the regularization loss of the generated masks.
@@ 5. Experiments
    Is the proposed parametric knowledge base model-agnostic? Does the proposed parametric knowledge base offer performance gain for all the base models?
@@
    Do the proposed two rules improve the quality of the parametric knowledge base?
@@
    Are the components of the parametric knowledge effective? How does the number of knowledge vectors per data instance influence the performance?
@@
    Are the objectives of the proposed constraints (i.e., essential & disentangled principles) achieved during training?
@@ 5.1.1. Datasets
  • Ali Display [^1] https://tianchi.aliyun.com/dataset/56. It is a dataset provided by Alibaba to estimate the click-through rate of Taobao display ads.
  • Eleme [^2] https://tianchi.aliyun.com/dataset/131047. It is a dataset provided by Eleme that offered take-away service to users.
The detailed descriptions for the used datasets are summarized in Table 1. For the two datasets, the clicked samples are treated as positive samples and the exposed but not clicked samples are treated as negative samples. We split the data using the global timestamps (Qin et al., 2020, 2021). For regular baseline models, we use the logs before $T_{1}$ for training, 50% of the logs after $T_{1}$ for validation, and 50% of the logs after $T_{1}$ for test. For those supplemented with extra knowledge vectors, we use the logs before $T_{0}$ for building the knowledge base, the logs between $[T_{0},T_{1})$ for training base models, 50% of the logs after $T_{1}$ for validation, and 50% of the logs after $T_{1}$ for test. ($T_{0}<T_{1}$). The training setup is illustrated in Figure 3.
Figure 3. The dataset splitting and training setup.
5.1.2. Baselines
Since our method is model-agnostic, we use different widely-used CTR models as our base models.
We first compared a group of feature interaction models. DeepFM (Guo et al., 2017) modeled 2-way pairwise feature interactions and used a DNN for generalization. xDeepFM (Lian et al., 2018) utilized the CIN network to model finite-order feature interactions. PNN (Qu et al., 2018) explicitly modeled the ”AND” relation between features by introducing the product layer. AutoInt (Song et al., 2019) built high-order feature interactions by the multi-head attention. DCN (Wang et al., 2017) learned both low-order feature interactions and high-order nonlinear features efficiently. We also compared with models that extracted multi-samples information for recommendation enhancement. DIN (Zhou et al., 2018) modeled user interests according to the attention scores between the candidate item and the target user’s historical clicked items.
5.1.3. Metrics
The evaluation metrics include area under ROC curve (AUC) and negative log-likehood (LogLoss).
5.1.4. Hyperparameters
5.2. Main Results (RQ1)
Table 2. Experiment results on AD and Eleme. Since EDK is model-agnostic, we test its compatibility on different backbones. N/A denotes the original recommendation backbone. Rel.Impr denotes the relative AUC improvement of EDK against each backbone model. The symbol * indicates statistically significant improvement with p-value $<0.001$.
@@ 5.2. Main Results (RQ1)
w/ EDK 0.6372* 0.1935* 0.6556* 0.0883*
@@
The overall performance of the proposed method is displayed in Table 2. Six widely used conventional recommendation models are used as the backbone models, the feature interaction operators of which include dnn-based, product-based, and attention-based methodologies. From the results, one can see that EDK consistently offers significant performance improvements for all the selected backbone models. The improvements are statistically significant under p-value $<0.001$. This validates that compressing knowledge of old data documents into parameters could preserve useful knowledge and offer performance gains for recommender systems. In addition, EDK shows superior model compatibility, which demonstrates the effectiveness of our established parametric knowledge base under the essential and disentangled constraints.
5.3. Ablation Study (RQ2-RQ3)
In this section, we conduct various ablation studies to validate the effectiveness of model designs.
5.3.1. Impact of the Two Proposed Principles (RQ2)

### v2

# FINED: Feed Instance-Wise Information Need with Essential and Disentangled Parametric Knowledge from the Past

@@ Abstract.
Recommender models play a vital role in various industrial scenarios, while often faced with the catastrophic forgetting problem caused by the fast shifting data distribution, e.g., the evolving user interests, click signals fluctuation during sales promotions, etc. To alleviate this problem, a common approach is to reuse knowledge from the historical data. However, preserving the vast and fast-accumulating data is hard, which causes dramatic storage overhead. Memorizing old data through a parametric knowledge base is then proposed, which compresses the vast amount of raw data into model parameters. Despite the flexibility, how to improve the memorization and generalization capabilities of the parametric knowledge base and suit the flexible information need of each instance are challenging. In this paper, we propose FINED to Feed INstance-wise information need with Essential and Disentangled parametric knowledge from past data for recommendation enhancement. Concretely, we train a knowledge extractor that extracts knowledge patterns of arbitrary order from past data and a knowledge encoder that memorizes the arbitrary order patterns, which serves as the retrieval key generator and memory network respectively in the following knowledge reusing phase. The whole process is regularized by the proposed two constraints, which improve the capabilities of the parametric knowledge base without increasing the size of it. The essential principle helps to compress the input into representative vectors that capture the task-relevant information and filter out the noisy information. The disentanglement principle reduces the redundancy of stored information and pushes the knowledge base to focus on capturing the disentangled invariant patterns. These two rules together promote rational compression of information for robust and generalized knowledge representations. Extensive experiments on two datasets justify the effectiveness of the proposed method.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: WWW; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Figure 1. (a) Overall process. (b) The essential principle for compression, which encourages a compressed representation that captures the fundamental knowledge of data and filters out the noises. (c) The disentangled principle for compression, which reduces the redundancy of stored patterns and decomposes the invariance for better generalization.
@@
To solve the challenges above, we propose the essential and disentangled parametric knowledge base, where two constraints regularize a parametric knowledge base for robust and generalized knowledge preserving without increasing the size of the knowledge base. Concretely, the Essential principle aims to capture the task-relevant information (the red part in Figure 1.(b)) and filter out the noisy information (the grey part in Figure 1.(b)) in the data, which follows the Information Bottleneck principle (Tishby et al., 2000a) that finds the optimal trade-off between compression and prediction. By striking a balance between these two aspects, the principle enables the model to preserve sufficient information of data while keeping the representation minimal and invariant (Achille and Soatto, 2018), which contributes to lower storage cost and higher generalization capability. The Disentanglement principle eliminates redundant information among different patterns (the grey part in Figure 1.(c)) and decomposes the invariance of different patterns. This reduction of redundancy helps to improve memorization within limited model parameters. And at the same time, as depicted in (Montero et al., 2020), improving disentanglement may contribute to the generalization. The two rules together contribute to compressed and rational knowledge vectors that capture the fundamental knowledge of past data and have good generalization ability for future distribution.
@@
  • We propose a parametric knowledge base that compresses the massive old data documents into robust and generalized parametric knowledge, which is a model-agnostic method that could offer diverse knowledge for various base models.
@@
  • The knowledge extraction and utilization of FINED is instance-level, which offers fine-grained information tailored for each instance for enhancement.
@@ 2. Related Work
2.1. Conventional Recommendation Backbones
Conventional recommenders capture collaborative signals based on feature interaction learning and user behavior modeling.
@@ 2.1. Conventional Recommendation Backbones
For feature interaction learning, there are plenty of works mining the interactive patterns of categorical data. FM (Rendle, 2010) captures feature interactions by inner products. FFM (Juan et al., 2016) further develops field-aware interactions by multiple embeddings settings. NeuFM (He et al., 2017) proposes to use deep neural networks, improving FM by incorporating a multi-layer perceptron. Wide & Deep (Cheng et al., 2016) combines the strengths of linear models and DNNs for memorization and generalization respectively. DeepFM (Guo et al., 2017) utilizes an FM layer to replace the wide component in (Cheng et al., 2016) to model the pairwise feature interactions. xDeepFM (Lian et al., 2018) introduces the cross layer to replace the wide component and constructs limited high order feature interactions. PNN (Qu et al., 2018) proposes a product layer to better model the “AND” relations between features. DCN (Wang et al., 2017) learns both low-dimensional feature crossing and high-dimensional nonlinear features efficiently. AutoInt (Song et al., 2019) learns high-order feature interactions via the self-attention mechanism.
@@
User behavior modeling plays a crucial role in recommendation systems by providing valuable insights into user preferences, interests, and behaviors. By understanding how users interact with a system and make choices, recommendation algorithms can effectively personalize and tailor recommendations to individual users. DIN (Zhou et al., 2018) models user interests with attention mechanism. It reduces the influence of historical behavior information of commodities that are not related to the currently estimated advertisement on the current click estimation judgment. DIEN (Zhou et al., 2019) further points out that user interests evolve over time and modeled the evolving interests of users with the GRU module. DSIN (Feng et al., 2019) splits different sessions and better models the evolving property of user interests. SIM (Pi et al., 2020) proposes to model long-term user behaviors and aggregates them using the fast SimHash algorithm.
2.2. Distribution Robust Recommenders
To deal with the fast-shifting distribution of recommendation datasets, various methodologies are came up, including continual learning methods, memory networks, and distributionally robust models.
Continual learning methods aim to continuously adapt a recommendation model to new data without forgetting previously learned knowledge, which need to adapt to users’ changing preferences, new items, or evolving contexts without the computational overhead of retraining from scratch. The main paradigms of continual learning can be summarized into regularization-based methods and methods using modular architecture. Regularization-based methods prevent the catastrophic forgetting of old data when the model learns from new data. They work by adding a regularization term to the loss function to constrain the updates on the model parameters (Lopez-Paz and Ranzato, 2017; Chaudhry et al., 2018; Mi et al., 2020; Wang et al., 2020), thus keeping them close to previous values unless warranted by strong evidence from new data. Methods that use modular architecture grow the base model when faced with newly arrived data (Riemer et al., 2018).
Memory networks in recommendation scenarios allow the system to store user interactions and context in a structured memory, enabling the model to refer to past interactions when generating recommendations. NMRN (Wang et al., 2018) proposes a streaming recommender with external memories. RUM (Chen et al., 2018) introduces a first-in-first-out writing mechanism, which stores and updates users’ historical records explicitly for better future prediction. KSR (Huang et al., 2018) integrates RNN-based networks with Key-Value Memory Networks and utilizes extensive knowledge base information to better encode user preferences. HPMN (Ren et al., 2019) propose a hierarchical periodic memory network model for lifelong sequential modeling in user response prediction, which can handle long-term sequential patterns and multi-scale sequential patterns of user interests. MIMN (Pi et al., 2019a) presents a co-design solution for handling long sequential user behavior data, combining a separate module called UIC (User Interest Center) in the serving system to manage user interests in real-time and a novel memory-based architecture to capture user interests from long sequential behavior data.
Distributionally Robust Optimization (DRO) in the context of recommendation systems is a powerful method aimed at improving the robustness and stability of recommendations under various types of data uncertainty and modeling assumptions. SML (Zhang et al., 2020) employs a neural network-based transfer component, enabling the retraining of recommender systems to focus on new data efficiently and effectively, thereby addressing overfitting and improving recommendation accuracy compared to full model retraining. InvCF (Zhang et al., 2023) generates unbiased preference representations resistant to changes in item popularity, demonstrating the ability to outperform leading models in terms of popularity generalization across multiple datasets and evaluation settings without requiring prior knowledge of popularity distributions.
3. Preliminaries
3.1. Problem Formulation
(1) $p(y|X,\theta),$
where $X=\left[x_{1},x_{2},\dots,x_{F}\right]$ is the input consisting of features from $F$ feature fields and $\theta$ is the model parameter. These methods make prediction based solely on target feature and model parameters. Another thread of methods that models user behavior patterns (Zhou et al., 2018, 2019; Pi et al., 2019b; Feng et al., 2019; Pi et al., 2020) shows important impacts on making personalized recommendation and brings great performance gains, which can be formulated as
(2) $p(y|X,D_{his},\theta),$
(3) $KB_{\theta*}\leftarrow D_{old},$
(4) $p(y|X,KB_{\theta*},\theta),$
(5) $p(y|X,D_{his},KB_{\theta*},\theta),$
@@ 3.2. Model Invariance & Generalization
Figure 2. The framework of FINED. In the knowledge compression stage, we compress the essential and disentangled knowledge within old data into the parametric knowledge base. Concretely, we extract patterns within data instances with the knowledge extractor and memorize them through the knowledge encoder that could deal with inputs of arbitrary scales. The overall knowledge compression process is regularized by two principles: essential and disentangled for better generalization and robustness. During prediction, the target could access the frozen knowledge base for instance-wise knowledge and adapt the knowledge to inject it into arbitrary recommendation backbone for enhanced prediction.
@@ 4.1.1. Essential & Disentangled Principles
  • Essential. $\min I(X;\{\mathbf{s_{j}}\})-\alpha I(\{\mathbf{s_{j}}\};y)$. The generated knowledge representations should contain as much of the task-relevant information of the original input and as less of task-irrelevant information. It helps to distill data into a sufficient and minimal representation, enabling better generalization and robustness.
@@ 4.1.4. Objective
where $l_{ce}$ denotes the cross entropy loss of predicting labels on $\mathbf{c}$, $l_{reg}$ corresponds to the two principles proposed in Equation (13), and $l_{0}$ denotes the regularization of the generated masks.
@@ 5. Experiments
    Can the proposed model beat the baseline models?
@@
    Is the proposed parametric knowledge base model-agnostic? Does FINED offer performance gains for different backbones?
@@
    Are the components of the parametric knowledge effective? Do the proposed constriants improve the quality of the parametric knowledge? How does the number of knowledge vectors per data instance influence the performance?
@@
    Are the knowledge patterns well captured? Are the objectives of the proposed constraints (i.e., essential & disentangled principles) achieved during training?
@@ 5.1. Experimental Setup
  • Ali Display [^1] https://tianchi.aliyun.com/dataset/56 is a dataset provided by Alibaba to estimate the click-through rate of Taobao display ads.
  • Eleme [^2] https://tianchi.aliyun.com/dataset/131047 is a dataset provided by Eleme that offered take-away service to users.
The detailed descriptions for the used datasets are summarized in Table 1. For the two datasets, the clicked samples are treated as positive samples and the exposed but not clicked samples are treated as negative samples. We split the data using the global timestamps (Qin et al., 2020, 2021). For regular baseline models, we use the logs before $T_{1}$ for training, 50% of the logs after $T_{1}$ for validation, and 50% of the logs after $T_{1}$ for test. For those supplemented with extra knowledge vectors, we use the logs before $T_{0}$ for building the knowledge base, the logs between $[T_{0},T_{1})$ for training base models, 50% of the logs after $T_{1}$ for validation, and 50% of the logs after $T_{1}$ for test. ($T_{0}<T_{1}$).
We compare the proposed FINED with four groups of baselines.
  • (Group 1) Memory networks. HPMN (Ren et al., 2019) and MIMN (Pi et al., 2019a) are memory networks that utilize external memory to preserve additional knowledge.
  • (Group 2) Incremental learning methods. ADER (Mi et al., 2020) and IncCTR (Wang et al., 2020) continually update the model parameters to adapt to different distributions.
  • (Group 3) Distributionally robust models. SML (Zhang et al., 2020) and InvCF (Zhang et al., 2023) are distributionally robust models that explore invariant patterns for shifting distributions.
  • (Group 4) Parametric knowledge base model. D2K (Qin et al., 2024) proposes to use model parameters learned on old data to construct knowledge base for future prediction.
Table 2. Comparisons with different baselines. Rel.Impr. denotes the relative AUC improvement of FINED against each baseline. The symbol * indicates statistically significant improvement with p-value $<0.001$.
Model AD Eleme
AUC Logloss Rel. Impr. AUC Logloss Rel. Impr.
Memory Network HPMN 0.6210 0.1970 2.11% 0.5984 0.0963 9.68%
MIMN 0.6178 0.1984 2.64% 0.5924 0.0987 10.79%
Incremental Learning ADER 0.6241 0.1944 2.20% 0.6122 0.0908 7.20%
IncCTR 0.6194 0.1955 2.97% 0.6033 0.0910 8.79%
Distributionally Robust Models SML 0.6221 0.1952 1.93% 0.5857 0.0927 12.05%
InvCF 0.6233 0.1945 2.33% 0.5961 0.0911 10.10%
Parametric Knowledge Base D2K 0.6341 0.1940 0.58% 0.6401 0.1343 2.53%
FINED 0.6378 0.1935 - 0.6563 0.0881 -
Table 3. Compatibility Study. N/A denotes the original recommendation backbone. Rel.Impr denotes the relative AUC improvement of FINED against each backbone model. The symbol * indicates statistically significant improvement with p-value $<0.001$.
@@
w/ FINED 0.6372* 0.1935* 0.6556* 0.0883*
Since our method is model-agnostic, we also test its compatibility with different backbone models. We evaluate the compatibility on six widely used conventional recommendation backbones, the feature interaction operators of which include dnn-based (xDeepFM (Lian et al., 2018), DCN (Wang et al., 2017)), product-based (DeepFM (Guo et al., 2017), PNN (Qu et al., 2018)), and attention-based (DIN (Zhou et al., 2018), AutoInt (Song et al., 2019)) methodologies.
The evaluation metrics include area under ROC curve (AUC) and negative log-likehood (LogLoss). For the hyperameters, details can be found in Section A in the Appendix.
5.2. Main Results (RQ1-RQ2)
The overall performance of the proposed method is displayed in Table 2. And the compatibility experiment on different backbones is displayed in Table 3.
@@ 5.2. Main Results (RQ1-RQ2)
(RQ1) For the model performance, from Table 2, one can see that FINED consistently outperform all the baseline models, including memory networks, incremental learning methods, distributionally robust models, and the recent proposed parametric knowledge base model. The improvements are statistically significant under p-value $<0.001$. This validates that compressing knowledge of old data documents into parameters could preserve useful knowledge and offer performance gains for recommender systems.
(RQ2) From Table 3, one can see that FINED consistently offers significant performance improvements for all the selected backbone models. The improvements are statistically significant under p-value $<0.001$. This shows that FINED has superior model compatibility, which demonstrates the effectiveness of our established parametric knowledge base under the essential and disentangled constraints.
5.3. Ablation Study (RQ3)
5.3.1. Impact of the Two Constraints